# SEGSRNet for Stereo-Endoscopic Image Super-Resolution and Surgical Instrument Segmentation

### v1

@@ Abstract
SEGSRNet addresses the challenge of precisely identifying surgical instruments in low-resolution stereo endoscopic images, a common issue in medical imaging and robotic surgery. Our innovative framework enhances image clarity and segmentation accuracy by applying state-of-the-art super-resolution techniques before segmentation. This ensures higher-quality inputs for more precise segmentation. SEGSRNet combines advanced feature extraction and attention mechanisms with spatial processing to sharpen image details, which is significant for accurate tool identification in medical images. Our proposed model outperforms current models including Dice, IoU, PSNR, and SSIM, SEGSRNet where it produces clearer and more accurate images for stereo endoscopic surgical imaging. SEGSRNet can provide image resolution and precise segmentation which can significantly enhance surgical accuracy and patient care outcomes.
@@ II Research Methodology
Our network tackles semantic segmentation in surgery by distinguishing instruments from the background in super-resolved images, enhancing both medical imaging and robotic surgery with an integrated super-resolution and segmentation framework.
@@ II-A1 Feature Extraction and Refinement Blocks
Our model features a Combined Channel and Spatial Attention Block (CCSB)[7], which includes a Channel Attention Block (CAB) for enhancing feature maps and a Spatial Channel Attention Block (SAB) for focusing on key regions. The features processed through CCSB are further refined using an Atrous Spatial Pyramid Pooling (ASPP) block and Residual Dense Blocks (RDBs). These components deepen feature extraction and create a comprehensive feature hierarchy, significantly improving the model’s performance in super-resolution and semantic segmentation tasks.
@@ II-C Datasets
Dataset 1 from the MICCAI 2018 Robotic Scene Segmentation Sub-Challenge [12] and Dataset 2 from the 2017 Robotic Instrument Segmentation Challenge [13] both have provided high-resolution ( pixels) stereo image pairs with camera calibrations from porcine training sessions. These datasets are pivotal for developing computational methods in robotic scene segmentation for endoscopic procedures.
@@ II-D Training Settings
The proposed model, developed in Pytorch 2.0 and trained on an Nvidia 3090Ti GPU, utilized Xavier initialization and the Adam optimizer with an initial learning rate of . It was trained over 100 epochs, with batch sizes 6 for scale two images and 5 for scale four.
@@ III Experimental Results
| Method | Scale | Dataset 1 | Dataset 2 |
@@ III-A Quantitative and Qualitative Results
The study measured the algorithm’s performance with PSNR and SSIM for image quality, and Dice Score and IoU for segmentation accuracy, with results in Tables I and II confirming its effectiveness.
For SR, Dataset 1 was used for training, and both datasets for testing, while segmentation tests on EndoVis 2017[13] highlighted our model’s superiority in binary and parts segmentation, outperforming models like U-Net [3] and TernausNet [4], as detailed in Table II and visually compared in Fig 4. However, our model had limitations in Type segmentation due to its emphasis on global contextual information, which is less suited for the fine-grained, pixel-level distinctions required among multiple complex classes, compared to the more generalized tasks in binary and parts segmentation.
Fig. 3 and Fig. 4 provide a comparative analysis of super-resolution performance and segmentation outcomes, respectively. Our model effectively computes disparities and corrects inaccuracies in SR, outperforming methods like Bicubic interpolation and DRRN, and demonstrates proficiency in binary, Parts, and Type segmentation tasks.
@@ IV Conclusion
SRSEGNet introduces a breakthrough in deep learning for super-resolution and segmentation in endoscopic vision, leveraging convolutional neural networks and SPP-LinkNet-34. Achieving high performance on the EndoVis 2017 dataset, SRSEGNet excels in binary segmentation with an IOU of 83.65% % and a Dice score of 89.80% %, effectively handling complex multi-class segmentation tasks.

### v2

@@ Abstract
SEGSRNet addresses the challenge of precisely identifying surgical instruments in low-resolution stereo endoscopic images, a common issue in medical imaging and robotic surgery. Our innovative framework enhances image clarity and segmentation accuracy by applying state-of-the-art super-resolution techniques before segmentation. This ensures higher-quality inputs for more precise segmentation. SEGSRNet combines advanced feature extraction and attention mechanisms with spatial processing to sharpen image details, which is significant for accurate tool identification in medical images. Our proposed model, SEGSRNet, surpasses existing models in evaluation metrics including PSNR and SSIM for super-resolution tasks, as well as IoU and Dice Score for segmentation. SEGSRNet can provide image resolution and precise segmentation which can significantly enhance surgical accuracy and patient care outcomes.
@@ II Research Methodology
Our proposed network addresses the challenge of semantic segmentation in surgery by distinguishing surgical instruments from the background in super-resolution images, enhancing both medical imaging and robotic surgery.
@@ II-A1 Feature Extraction and Refinement Blocks
Our model features a Combined Channel and Spatial Attention Block (CCSB)[7], which includes a Channel Attention Block (CAB) for enhancing feature maps and a Spatial Channel Attention Block (SAB) for focusing on key regions. The features processed through CCSB are further refined using an Atrous Spatial Pyramid Pooling (ASPP) block and Residual Dense Blocks (RDBs). These components deepen feature extraction and create a comprehensive feature hierarchy, significantly improving the model’s performance in super-resolution.
@@ II-C Datasets
We use The MICCAI 2018 Robotic Scene Segmentation Sub-Challenge [12] (”MICCAI 2018”) and the MICCAI 2017 Robotic Instrument Segmentation Challenge [13] (”EndoVis 2017”) are the datasets for the robotic scene segmentation in endoscopic procedures. Both challenges provide high-resolution stereo image pairs pixels) captured during porcine training sessions, along with camera calibrations. For the SR task, we trained our models using the MICCAI 2018 dataset and then evaluated their performance using both the MICCAI 2018 and EndoVis 2017 datasets. Conversely, for segmentation, we conducted both training and testing exclusively on the EndoVis 2017 dataset. We calculate Peak signal-to-noise ratio (PSNR) and Structural Similarity (SSIM) as evaluation metrics. For segmentation performace, we use 10-folds cross-validation to measure segmentation Intersection over Union (IoU) on EndoVis 2017 and calculate the mean and standard deviation (STD) of the validation set.
@@ II-D Training Settings
Our proposed model was implemented using the Pytorch 2.0 framework and trained on a Nvidia 3090Ti GPU. We employed Xavier initialization for the model’s parameters and used Adam optimizer with an initial learning rate of . All models are trained for 100 epochs. We used a batch size of 6 for images at scale two () and a batch size of 5 for images at scale four ().
@@ III Experimental Results
| Method | Scale |
MICCAI 2018
(PSNR/SSIM) |
EndoVis 2017
(PSNR/SSIM) |
@@
| Method | Scale |
MICCAI 2018
(PSNR/SSIM) |
EndoVis 2017
(PSNR/SSIM) |
@@ III-A Quantitative and Qualitative Results
Our model outperforms traditional U-Net by 9.81% and 27.60% in binary and parts segmentation in terms of IoU. However, our model had limitations in type segmentation due to its emphasis on global contextual information which is less suited for the fine-grained, pixel-level distinctions required among multiple complex classes, compared to the more generalized tasks in binary and parts segmentation.
The super-resolution results show that SEGSRNet corrects inaccuracies and computes disparities more effectively than traditional methods such as Bicubic interpolation and DRRN (Fig 3). Overall, SEGSRNet outperforms traditional model in both and . After applying SR, we illustrates the model’s proficiency in various segmentation tasks, including binary, Parts, and Type segmentation, highlighting its superior performance in accurately segmenting different image components (Fig 4).
@@ IV Conclusion
SRSEGNet introduces a breakthrough in deep learning for super-resolution and segmentation in endoscopic vision, leveraging convolutional neural networks and SPP-LinkNet-34. Achieving high performance on the EndoVis 2017 dataset, SRSEGNet excels in binary segmentation with an IoU of 83.65% and a Dice score of 89.80%, effectively handling complex multi-class segmentation tasks.